OpenStudy (anonymous):

pythagorean theorem: C = √a^2 + √b^2 Should c = √a^2 + √b^2 have a +- symbol in front of the c? Why or why not? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

the pythagorean theorem is a^2 + b^2 = c^2

jimthompson5910 (jim_thompson5910):

if you apply square roots to both sides, you get \[\Large a^2 + b^2 = c^2\] \[\Large \sqrt{a^2 + b^2} = \sqrt{c^2}\] notice how there is one root on the left side and it covers ALL of the left side

jimthompson5910 (jim_thompson5910):

on the right side, the square and square root cancel, leaving just c behind

jimthompson5910 (jim_thompson5910):

so that's why \[\Large c = \sqrt{a^2 + b^2}\]

jimthompson5910 (jim_thompson5910):

and c is some distance c is never negative so you don't have to worry about the plus/minus here. just focus on the plus

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

there won't be a plus/minus symbol

jimthompson5910 (jim_thompson5910):

there shouldn't be one

jimthompson5910 (jim_thompson5910):

it's simply \[\Large c = \sqrt{a^2 + b^2}\]

jimthompson5910 (jim_thompson5910):

and it is NOT this \[\Large c = \pm\sqrt{a^2 + b^2}\] the minus implies that we have negative distance which doesn't make sense

OpenStudy (anonymous):